OpenStudy (anonymous):

Jennie is working at a lighthouse. From the top of the lighthouse, she spots a boat at an angle of depression of 25(degrees) She knows that the landmark rock located next to the boat is 600 feet from the bottom of the lighthouse. What is Jennie’s elevation? Round to the nearest foot.

OpenStudy (anonymous):

Can you draw this situation? Try.

OpenStudy (anonymous):

That was the first thing that I did.

OpenStudy (anonymous):

can you just help me I have a quiz supposed to be done in five minutes I don't have time?//

OpenStudy (anonymous):

The 25 degrees is measure from the horizontal _down_ to Jennie's line of sight to the boat. So the angle between her line of sight and the lighthouse is 90 - 25 = 65 degrees. The ratio of 600 to the height of the lighthouse is the tangent of this 65 degrees. Calling h the required height we have : tan 65 = 600 / h or h tan 65 = 600 or h = 600 / tan 65, which is about 279.8 feet

OpenStudy (anonymous):

so the answer is?

OpenStudy (anonymous):

280 feet